# thanks: Equal contributionthanks: Equal contribution

### v1

@@ V.2.1 Comparison across infrastructures
Our first suite of tests consists of evolving the same TOV initial data across different infrastructures, comparing the following codes: origIGM and IllinoisGRMHD within the Einstein Toolkit, and NRPyCartMHD. The initial central density of the star is $\rho_{c}=1.28\times 10^{-3}$ in code units, and we adopt an initially cold hybrid EoS with $K_{0}=100$ and $\Gamma_{0}=\Gamma_{\rm th}=2$. The TOV initial data do not include magnetic fields, making this a pure hydrodynamics test. The spacetime is evolved using Baikal [Ruchlin:2018, Etienne_nrpytutorial], available within both the Einstein Toolkit and BlackHoles@Home.
Baikal evolves the spacetime variables using the BSSN formulation [Brown_2009], and employs the $1+\log$ lapse [Campanelli_2006] and a second-order, non-covariant, advective shift evolution [Meter_2006] as gauge conditions, i.e.
@@
where $K$ is the trace of the extrinsic curvature $K_{ij}$, $\bar{\Lambda}^{i}$ and $B^{i}$ are auxiliary variables, and $\eta$ is the shift damping parameter. We set $\eta=1/3$, use Kreiss–Oliger dissipation at a strength of $0.2$ [Kreiss_1973], and use fourth-order finite differencing for the spacetime fields. For the hydrodynamic fields we use PPM reconstruction, the Noble2D conservative-to-primitive solver, and set the atmospheric density to $\rho_{\mathrm{atm}}=\rho_{c}\times 10^{-9}$. In all tests we ensure that the diameter of the star is sampled with at least ${\approx}30$ points along a coordinate axis. These codes evolve the initial data forward in time using the method of lines and an RK4 scheme and, apart from NRPyCartMHD, use five levels of fixed mesh refinement (FMR) in Cartesian coordinates with the outer boundary placed at ${\pm}240$ in all three directions. We evolve the initial data at two resolutions, MR and HR, corresponding to resolutions on the coarsest level of $dx=8.0$ and $dx=5.0$, respectively. Since NRPyCartMHD supports only a single Cartesian grid, we place its outer boundary at ${\pm}40$ in all three directions and choose the grid spacing to match that of the finest refinement level of the FMR hierarchy.
In the top panel of Fig. 4 we show results of the time evolution of the normalized change in central density for all codes, truncating the results from NRPyCartMHD when outer boundary effects become apparent, with all the GRHayL-based codes using the full suite of changes from Sec. III. These results demonstrate that the frequencies of the fundamental modes of oscillation for the star are recovered [Font_2000_nonlinear_tests, Font_2002]. Further, the amplitude of these oscillations, which are driven by discretization errors in the nominally static initial data, is expected to converge to zero with increasing resolution. In the bottom panel of Fig. 4 we plot the approximate convergence order of the error in central density for origIGM and IllinoisGRMHD. We observe that they have a similar convergence order at around $n=2.5$, consistent with what was found in previous studies (e.g., [Werneck_IGM]).
Figure 4: Top: Comparison of non-magnetized TOV initial data evolution between origIGM, IllinoisGRMHD, GRHayLHDX, and NRPyCartMHD, each at two resolutions. IllinoisGRMHD and NRPyCartMHD are equivalent until outer boundary effects are apparent, therefore we truncate the results of NRPyCartMHD before they manifest. Bottom: Comparison of convergence order between origIGM and IllinoisGRMHD. IllinoisGRMHD consistently outperforms origIGM.
Figure 5: Comparison of non-magnetized TOV initial data evolution between origIGM and various changes made to IllinoisGRMHD during development.
@@
Figure 6: Relative difference between IllinoisGRMHD and NRPyCartMHD, comparing two outer boundary locations while keeping the resolution fixed and using a single refinement level.
@@ V.3 Piecewise polytrope EoS study
Figure 8: Fourier transform of time series of the central density evolution for a TOV in dynamical spacetime with the piecewise polytropic representation of the SLy EoS. Vertical dashed lines indicate the frequencies of the lowest four radial oscillation modes for the SLy EoS [Sen:2022kva].
V.4 BNS Inspiral, Merger, and Collapse of HMNS
@@ V.4.1 Hybrid EoS, Simple Polytrope
Figure 10: Comparison of the maximum density for origIGM with randomly perturbed initial data from hypermassive neutron star formation to collapse.
@@
We use LORENE initial data for an equal-mass BNS system with an initial separation of $45$ km. Each star has mass $M\approx 1.625\,M_{\odot}$ and radius $R\approx 13.67\,\mathrm{km}$. We use the Seed_Magnetic_Fields thorn to set up the initial magnetic fields and the Baikal thorn to evolve the spacetime. Further details regarding the AMR grids and thorn setup are provided in the hybrid BNS parameter file packaged with the IllinoisGRMHD thorn.
Due to the various improvements (detailed in Sec. III), the codes no longer agree at the round-off-level. To assess their level of agreement, we perform three BNS simulations with both evolution codes to analyze their behavior. One is the control with no perturbations, and the other two perturb the initial $\bm{P}$ data at the 9th and 14th significant digits, respectively.
Fig. 9 shows the quantity $\rho_{\mathrm{max}}/\rho_{0}$ during the inspiral phase of the simulation for all six simulations. The two codes show the same behavior, and the differences follow the same trends as the TOV simulations. At this stage of the simulation, the perturbations show no significant effect on the central density of the stars. The central densities dissipate away faster in IllinoisGRMHD due to the change to the $P_{\mathrm{min}}$ floor, which has a slightly stronger effect in these simulations than in the TOV simulations.
The perturbative effects start significantly affecting the evolution during the hypermassive neutron star phase, where even small changes can affect the evolution of the system. Fig. 10 compares $\rho_{\mathrm{max}}$ from the origIGM simulations (unperturbed and perturbed at the 14th digit) with that from the unperturbed IllinoisGRMHD simulation. We cut off the data once the simulation reaches the maximum allowed density for the simulation. As shown, perturbations to initial data cause the number of oscillations to change in the original origIGM code. The new IllinoisGRMHD has round-off-level differences at every Con2Prim evaluation relative to the original code due to code improvements and changes, which would similarly cause differences between the old and new results. Despite these cumulative perturbations, IllinoisGRMHD’s results are still on the order of a round-off-level perturbation relative to the original code. Given the unstable nature of the hypermassive neutron star and the repeated perturbations caused by various code improvements and simplifications, IllinoisGRMHD displays remarkable agreement with origIGM.
@@
(e) $t=3600$ (post-collapse).
(f) $t=4640$ (post-collapse).
Figure 12: Same as Fig. 11 for the post-merger phase.
@@
In Fig. 12, we show three spatial slices for the post-merger. Note that the collapse time is different between the codes. The first two sets of plots are for the same time and display very similar behavior. The magnetic field strength again follows the matter flows better in IllinoisGRMHD, with the matter tails coming off the merging binary having accompanying magnetic components that are missing in origIGM. Fig. 12e and Fig. 12f show the post-collapse system after the central mass has settled down into a black hole with inspiraling matter. Both codes show very similar dynamics and remain stable out to $t=6000$.
To illustrate the differences more clearly, Fig. 13 and Fig. 14 show the density and magnetic field strength for both codes. origIGM has very noticeable distortions in magnetic field topology near the star surface, while IllinoisGRMHD has a much smoother profile. Further, the magnetic field in IllinoisGRMHD follows the distortion of the stars, with the deformation leading to merger causing a deformation of the magnetic field strength, as seen in Fig. 13. In contrast, origIGM remains nearly spherical.
As the stars enter the merger, the inspiral rate outpaces the magnetic field’s ability to keep up, leading to the bulge in magnetic field strength visible in Fig. 14. The deformation towards positive $x$ for the lower star is the same deformation present in Fig. 13, as this feature fails to follow the star’s rotation through the last quarter-rotation into the merger. This is not apparent in origIGM since it remains nearly spherical all the way to merger.
@@ V.4.2 Hybrid EoS, Piecewise Polytrope
We set up a grid with eight refinement levels by factors of two, with the resolution at the finest level $\Delta x_{8}\approx 185\penalty 10000\ \mathrm{m}$. After BH formation, two additional refinement levels are added to resolve the puncture, making the highest grid resolution $\Delta x_{10}\approx 46\penalty 10000\ \mathrm{m}$.
@@ VI Concluding remarks
GRHayL provides an effective method for quickly developing GRMHD evolution code in new infrastructures. Any changes to the library are validated by continuous integration testing (see Appendix B). Few GRMHD codes perform regular testing at the level of individual functions (and demand round-off-level agreement), and these tests represent professional industry standards that are rarely implemented in open-source scientific codes.
The current library contains all the necessary pieces to quickly implement an IllinoisGRMHD-like code in new infrastructures. This includes multiple EoS types, Con2Prim and reconstruction methods, hydrodynamic and induction equation RHSs, and neutrino leakage. In addition to the current features, plans include adding more advanced radiation transport schemes, more atmosphere prescriptions and more Con2Prim options in the near future. Work is already underway to extend the Noble implementations to support tabulated EoSs, as well as adding the 2D and 3D Cerdá-Durán et al. routines [Cerda_2008]. Support for M1 closure neutrino transport based on [Radice:2021jtw] is also planned.
Implementation of BlackHoles@Home’s multi-patch, multi-coordinate infrastructure within GRoovy will enable highly efficient BNS simulations. Work also continues towards adding magnetic evolution to the curvilinear code, which must be handled with particular care due to the combination of staggered grids and coordinate singularities.
@@ Acknowledgements
T. Pierre Jacques gratefully acknowledges funding support from NASA FINESST-80NSSC23K1437, West Virginia University’s Chancellor’s Scholarship program, and the Southern Regional Education Board’s Dissertation Award program. Z. B. Etienne gratefully acknowledges funding support from NSF awards AST-2227080, PHY-2110352, OAC-2227105, and NASA awards ISFM-80NSSC21K1179 and TCAN-80NSSC18K1488.
@@ Appendix A Basic Equations
where $P$ is the pressure, $v^{i}\equiv u^{i}/u^{0}$ is the fluid three-velocity, and $S$ is the entropy. Note that this velocity is different from the velocity used in the Valencia formulation [Banyuls_1997, Marti_1991]
@@
Similarly, there are two common definitions of $B^{\mu}$, the magnetic field measured by Eulerian observers. The first, adopted by e.g., [Duez_IGM], is
@@
while the second, adopted by e.g., [GRHydro], is
@@ Appendix B Continuous Integration Testing
The industry standard for code testing is to have regularly scheduled tests which validate individual components, or units. These unit tests validate each individual component so that if a change breaks the code, it is clear which piece has diverged from the expected behavior. Since many codebases are worked on by large groups of people who may be implementing many different features simultaneously, it is standard to use a technique known as continuous integration (CI) testing. CI runs unit tests every time the triggers are met, such as whenever someone pushes to the repository. The library is also indirectly tested via the Einstein Toolkit’s CI, as several tests now include GRHayL-based thorns. However, we only discuss testing performed by the library and not third parties.
@@ B.1 Code Test Methodology
GRHayL has been extensively tested, and automated tests immediately validate any changes to the code. The tests read input data $u_{\mathrm{in}}$ and compute an output $v_{\mathrm{new}}$. We compare this against a trusted reference output $v_{\mathrm{old}}$. To estimate round-off-level sensitivity, we additionally compute $v_{\mathrm{pert}}$ by perturbing $u_{\mathrm{in}}$ and re-running the trusted code.
@@
As previously mentioned, the data used for these comparisons is kept in the GRHayL_TestData repository. This serves to keep GRHayL as small as possible and also permits more and larger tests since we do not have to be as worried about the file sizes. These tests are optionally compiled along with the library, and the source code is in the Unit_Tests directory. Any test that requires data from the GRHayL_TestData repo also has a data generation file in the data_gen subdirectory.
The tests can be separated into two categories. The tests prefaced with ET_Legacy have matching code in origIGM. As such, the only data generated by GRHayL is the input data. A separate branch of origIGM (GRHayLTestPatch) contains the code changes and a parfile to read in this data. The branched origIGM code produces the trusted outputs $v_{\mathrm{old}}$ and $v_{\mathrm{pert}}$ as binary files for all the ET_Legacy tests (which are also stored in the GRHayL_TestData repository). The GRHayL library computes $v_{\mathrm{new}}$ from the same input data, and these values are compared to $v_{\mathrm{old}}$ and $v_{\mathrm{pert}}$, ensuring that the GRHayL code matches the progenitor code in origIGM. For new code, the data and data-generating code are grouped into directories by gem. The data-generating code for functions new to GRHayL produces input, output, and perturbed output binary files.

### v2

@@ V.2.1 Comparison across infrastructures
Our first suite of tests consists of evolving the same TOV initial data across different infrastructures, comparing the following codes: origIGM and IllinoisGRMHD within the Einstein Toolkit, and NRPyCartMHD. The initial central density of the star is $\rho_{c}=1.28\times 10^{-3}$ in code units, and we adopt an initially cold hybrid EoS with $K_{0}=100$ and $\Gamma_{0}=\Gamma_{\rm th}=2$. The TOV initial data do not include magnetic fields, making this a pure hydrodynamics test. The spacetime is evolved using Baikal [100, 46], available within both the Einstein Toolkit and BlackHoles@Home.
Baikal evolves the spacetime variables using the BSSN formulation [25], and employs the $1+\log$ lapse [27] and a second-order, noncovariant, advective shift evolution [113] as gauge conditions, i.e.
@@
where $K$ is the trace of the extrinsic curvature $K_{ij}$, $\bar{\Lambda}^{i}$ and $B^{i}$ are auxiliary variables, and $\eta$ is the shift damping parameter. We set $\eta=1/3$, use Kreiss-Oliger dissipation at a strength of $0.2$ [69], and use fourth-order finite differencing for the spacetime fields. For the hydrodynamic fields we use PPM reconstruction, the Noble2D conservative-to-primitive solver, and set the atmospheric density to $\rho_{\mathrm{atm}}=\rho_{c}\times 10^{-9}$. In all tests we ensure that the diameter of the star is sampled with at least ${\approx}30$ points along a coordinate axis. These codes evolve the initial data forward in time using the method of lines and a fourth-order Runge-Kutta scheme and, apart from NRPyCartMHD, use five levels of fixed mesh refinement (FMR) in Cartesian coordinates with the outer boundary placed at ${\pm}240$ in all three directions. We evolve the initial data at two resolutions, MR and HR, corresponding to resolutions on the coarsest level of $dx=8.0$ and $dx=5.0$, respectively. Since NRPyCartMHD supports only a single Cartesian grid, we place its outer boundary at ${\pm}40$ in all three directions and choose the grid spacing to match that of the finest refinement level of the FMR hierarchy.
In the top panel of Fig. 4 we show results of the time evolution of the normalized change in central density for all codes, truncating the results from NRPyCartMHD when outer boundary effects become apparent, with all the GRHayL-based codes using the full suite of changes from Sec. III. These results demonstrate that the frequencies of the fundamental modes of oscillation for the star are recovered [50, 48]. Further, the amplitude of these oscillations, which are driven by discretization errors in the nominally static initial data, is expected to converge to zero with increasing resolution. In the bottom panel of Fig. 4 we plot the approximate convergence order of the error in central density for origIGM and IllinoisGRMHD. We observe that they have a similar convergence order at around $n=2.5$, consistent with what was found in previous studies (e.g., [115]).
Figure 4: Top: comparison of nonmagnetized TOV initial data evolution between origIGM, IllinoisGRMHD, GRHayLHDX, and NRPyCartMHD, each at two resolutions. IllinoisGRMHD and NRPyCartMHD are equivalent until outer boundary effects are apparent, therefore we truncate the results of NRPyCartMHD before they manifest. Bottom: comparison of convergence order between origIGM and IllinoisGRMHD. IllinoisGRMHD consistently outperforms origIGM.
Figure 5: Comparison of nonmagnetized TOV initial data evolution between origIGM and various changes made to IllinoisGRMHD during development.
@@
Figure 6: Relative difference between IllinoisGRMHD and NRPyCartMHD, comparing two outer boundary (OB) locations while keeping the resolution fixed and using a single refinement level.
@@ V.3 Piecewise polytrope EoS study
Figure 8: Fourier transform of time series of the central density evolution for a TOV in dynamical spacetime with the piecewise polytropic representation of the SLy EoS. Vertical dashed lines indicate the frequencies of the lowest four radial oscillation modes for the SLy EoS [106].
V.4 BNS Inspiral, Merger, and Collapse of Hypermassive Neutron Star
@@ V.4.1 Hybrid EoS, Simple Polytrope
Figure 10: Comparison of the maximum density for origIGM with randomly perturbed initial data from hypermassive neutron star formation to collapse. We note that the near overlap between IllinoisGRMHD and the perturbed run with origIGM is purely coincidental.
@@
[…]
Because of the various improvements (detailed in Sec. III), the codes no longer agree at the round-off level. To assess their level of agreement, we perform three BNS simulations with both evolution codes to analyze their behavior. One is the control with no perturbations, and the other two perturb the initial $\bm{P}$ data at the 9th and 14th significant digits, respectively.
Figure 9 shows the quantity $\rho_{\mathrm{max}}/\rho_{0}$ during the inspiral phase of the simulation for all six simulations. The two codes show the same behavior, and the differences follow the same trends as the TOV simulations. At this stage of the simulation, the perturbations show no significant effect on the central density of the stars. The central densities dissipate away faster in IllinoisGRMHD due to the change to the $P_{\mathrm{min}}$ floor, which has a slightly stronger effect in these simulations than in the TOV simulations.
The perturbative effects start significantly affecting the evolution during the hypermassive neutron star phase, where even small changes can affect the evolution of the system. Figure 10 compares $\rho_{\mathrm{max}}$ between origIGM simulations (unperturbed and perturbed at the 14th digit) and an unperturbed IllinoisGRMHD simulation. We cut off the data once the simulation reaches the maximum allowed density for the simulation. As shown, perturbations to initial data cause the number of oscillations to change in the original origIGM code. The new IllinoisGRMHD has round-off-level differences at every Con2Prim evaluation relative to the original code due to code improvements and changes, which would similarly cause differences between the old and new results. Despite these cumulative perturbations, IllinoisGRMHD’s results are still on the order of a round-off-level perturbation relative to the original code. Given the unstable nature of the hypermassive neutron star and the repeated perturbations caused by various code improvements and simplifications, IllinoisGRMHD displays remarkable agreement with origIGM.
@@
(e) $t=3600$ (postcollapse).
(f) $t=4640$ (postcollapse).
Figure 12: Same as Fig. 11 for the postmerger phase.
@@
In Fig. 12, we show three spatial slices for the postmerger. Note that the collapse time is different between the codes. The first two sets of plots are for the same time and display very similar behavior. The magnetic field strength again follows the matter flows better in IllinoisGRMHD, with the matter tails coming off the merging binary having accompanying magnetic components that are missing in origIGM. Figures 12e and 12f show the postcollapse system after the central mass has settled down into a black hole with inspiraling matter. Both codes show very similar dynamics and remain stable out to $t=6000$.
To illustrate the differences more clearly, Figures 13 and 14 show the density and magnetic field strength for both codes. origIGM has very noticeable distortions in magnetic field topology near the star surface, while IllinoisGRMHD has a much smoother profile. Further, the magnetic field in IllinoisGRMHD follows the distortion of the stars, with the deformation leading to merger causing a deformation of the magnetic field strength, as seen in Fig. 13. In contrast, origIGM remains nearly spherical.
As the stars enter the merger, the inspiral rate outpaces the magnetic field’s ability to keep up, leading to the bulge in magnetic field strength visible in Fig. 14. The deformation toward positive $x$ for the lower star is the same deformation present in Fig. 13, as this feature fails to follow the star’s rotation through the last quarter rotation into the merger. This is not apparent in origIGM since it remains nearly spherical all the way to merger.
@@ V.4.2 Hybrid EoS, Piecewise Polytrope
We set up a grid with eight refinement levels by factors of 2, with the resolution at the finest level $\Delta x_{8}\approx 185\penalty 10000\ \mathrm{m}$. After black hole (BH) formation, two additional refinement levels are added to resolve the puncture, making the highest grid resolution $\Delta x_{10}\approx 46\penalty 10000\ \mathrm{m}$.
@@ VI Concluding remarks
GRHayL provides an effective method for quickly developing GRMHD evolution code in new infrastructures. Any changes to the library are validated by continuous-integration testing (see Appendix B). Few GRMHD codes perform regular testing at the level of individual functions (and demand round-off-level agreement), and these tests represent professional industry standards that are rarely implemented in open-source scientific codes.
The current library contains all the necessary pieces to quickly implement an IllinoisGRMHD-like code in new infrastructures. This includes multiple EoS types, Con2Prim and reconstruction methods, hydrodynamic and induction equation RHSs, and neutrino leakage. In addition to the current features, plans include adding more advanced radiation transport schemes, more atmosphere prescriptions and more Con2Prim options in the near future. Work is already underway to extend the Noble implementations to support tabulated EoS, as well as adding the 2D and 3D Cerdá-Durán et al. routines [28]. Support for M1 closure neutrino transport based on [94] is also planned.
Implementation of BlackHoles@Home’s multipatch, multicoordinate infrastructure within GRoovy will enable highly efficient BNS simulations. Work also continues toward adding magnetic evolution to the curvilinear code, which must be handled with particular care due to the combination of staggered grids and coordinate singularities.
@@ Acknowledgments
T.P.J. gratefully acknowledges funding support from NASA FINESST-80NSSC23K1437, West Virginia University’s Chancellor’s Scholarship program, and the Southern Regional Education Board’s Dissertation Award program. Z.B.E. gratefully acknowledges funding support from NSF Awards No. AST-2227080, PHY-2110352, OAC-2227105, and NASA Awards No. ISFM-80NSSC21K1179 and TCAN-80NSSC18K1488.
@@
Data Availability
The data that support the findings of this article are not publicly available upon publication because it is not technically feasible and/or the cost of preparing, depositing, and hosting the data would be prohibitive within the terms of this research project. The data are available from the authors upon reasonable request.
@@ Appendix A Basic Equations
where $P$ is the pressure, $v^{i}\equiv u^{i}/u^{0}$ is the fluid three-velocity, and $S$ is the entropy. Note that this velocity is different from the velocity used in the Valencia formulation [23, 73]
@@
Similarly, there are two common definitions of $B^{\mu}$, the magnetic field measured by Eulerian observers. The first, adopted by e.g., [38], is
@@
while the second, adopted by e.g., [82], is
@@ Appendix B Continuous Integration Testing
The industry standard for code testing is to have regularly scheduled tests which validate individual components, or units. These unit tests validate each individual component so that if a change breaks the code, it is clear which piece has diverged from the expected behavior. Since many codebases are worked on by large groups of people who may be implementing many different features simultaneously, it is standard to use a technique known as continuous-integration (CI) testing. CI runs unit tests every time the triggers are met, such as whenever someone pushes to the repository. The library is also indirectly tested via the Einstein Toolkit’s CI, as several tests now include GRHayL-based thorns. However, we only discuss testing performed by the library and not third parties.
@@ B.1 Code Test Methodology
GRHayL has been extensively tested, and automated tests immediately validate any changes to the code. The tests read input data $u_{\mathrm{in}}$ and compute an output $v_{\mathrm{new}}$. We compare this against a trusted reference output $v_{\mathrm{old}}$. To estimate round-off-level sensitivity, we additionally compute $v_{\mathrm{pert}}$ by perturbing $u_{\mathrm{in}}$ and rerunning the trusted code.
@@
As previously mentioned, the data used for these comparisons are kept in the GRHayL_TestData repository. This serves to keep GRHayL as small as possible and also permits more and larger tests since we do not have to be as worried about the file sizes. These tests are optionally compiled along with the library, and the source code is in the Unit_Tests directory. Any test that requires data from the GRHayL_TestData repository also has a data generation file in the data_gen subdirectory.
The tests can be separated into two categories. The tests prefaced with ET_Legacy have matching code in origIGM. As such, the only data generated by GRHayL are the input data. A separate branch of origIGM (GRHayLTestPatch) contains the code changes and a parfile to read in this data. The branched origIGM code produces the trusted outputs $v_{\mathrm{old}}$ and $v_{\mathrm{pert}}$ as binary files for all the ET_Legacy tests (which are also stored in the GRHayL_TestData repository). The GRHayL library computes $v_{\mathrm{new}}$ from the same input data, and these values are compared to $v_{\mathrm{old}}$ and $v_{\mathrm{pert}}$, ensuring that the GRHayL code matches the progenitor code in origIGM. For new code, the data and data-generating code are grouped into directories by gem. The data-generating code for functions new to GRHayL produces input, output, and perturbed output binary files.